OpenStudy (anonymous):

Fiona has proved that a function, f(x), is an arithmetic sequence. How did she prove that? A. She showed that an explicit formula could be created. B. She showed that a recursive formula could be created. C. She showed that f(n) ÷ f(n - 1) was a constant ratio. D. She showed that f(n) - f(n - 1) was a constant difference.

OpenStudy (triciaal):

arithmetic sequence first term a next term a + d where d is the difference between the 2 consecutive terms nth term a + (n-1)d

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ex.: 1 3 5 7 ... 3 - 1 = 5 - 3 = 7 - 5 = ;.... = 2

OpenStudy (anonymous):

thnx :)